# Time Travel in LLMs: Tracing Data Contamination in Large Language Models

### v1

@@ 3.2 Detecting Partition-level Contamination
To generalize from instance-level contamination scores to partition-level discrete decisions (i.e., the partition is/is not contaminated) we take advantage of two observations:
@@
Idea 1: A dataset is likely to be contaminated if the average overlap score (as measured by ROUGE or BLEURT) observed with the guided instruction is significantly larger than the one measured with the general instruction. The motivation behind this idea is that, since the only difference between the two instructions is that the guided instruction contains the dataset and partition names, the improvement can only be explained through contamination.
Idea 2: A dataset is likely to be contaminated if GPT-4 detects at least one exact match or at least two near-exact matches. The intuition behind this idea is that even a small contaminated part of the sample of instances is likely to be indicative of a larger dataset leak.
We propose two algorithms based on these ideas, with each algorithm implementing one of the ideas respectively.
@@
Algorithm 1: A dataset partition is labeled as contaminated if the overlap score (as provided by BLEURT or ROUGE) with the guided instruction on a sample of $k=10$ instances is significantly better than the score with the general instruction under a non-parametric bootstrap resampling test. [^2] Details of our bootstrap resampling method can be found in Appendix B.
@@
The advantage of this algorithm is that it is non-parametric, i.e., we do not need to decide on an arbitrary threshold on the ROUGE or BLEURT scores to indicate contamination. However, its drawback is that even a significant increase in overlap may still come from generated instances that a human would not consider an exact or near-exact match. Algorithm 2 addresses this limitation.
@@
Algorithm 2: A dataset partition is labeled as contaminated if GPT-4 labels at least one instance as an exact match or at least two instances as near-exact matches in a sample of $k=10$ instances. All the instances in this sample are generated using the guided instruction.
We evaluate both these algorithms next.
@@ 4 Experimental Setup
Data: Our study employs seven datasets derived from various tasks, namely classification, summarization, and NLI. The datasets in question involve IMDB (Maas et al. 2011), AG News (Zhang, Zhao, and LeCun 2015), Yelp Full Reviews (Zhang, Zhao, and LeCun 2015), SAMSum (Gliwa et al. 2019), XSum (Narayan, Cohen, and Lapata 2018), WNLI (Wang et al. 2018), and RTE (Wang et al. 2019). In order to ensure a comprehensive experimental setup, all our experiments are carried out on both the training and test/validation splits of the aforesaid datasets. We make use of the publicly available divisions, working with the training and test splits for each. However, for the last two datasets, only the validation splits were publicly accessible with their labels. Considering our research’s emphasis on pinpointing data contamination with minimal dataset instances, the resource constraints, and our intention to facilitate the replication of this approach by other researchers, we randomly chose 10 instances from each split for our experiments.
@@
Setting: We use snapshots of GPT-3.5 and GPT-4 from June 13, 2023—specifically gpt-3.5-turbo-0613 and gpt-4-0613—both accessed via the OpenAI API, as our foundational LLMs. To obtain deterministic results, we set the temperature to zero, and capped the maximum completion length at 500 tokens. Contrarily, our comparative method uses the chat user interface (UI), which we also leveraged for conducting the experiment under this method. Specifically, we used the UI versions of GPT-4 and GPT-3.5 that were released on July 20, 2023.
@@
Human Evaluation: We undertake a human evaluation, led by a domain expert, to characterize contamination by identifying both exact matches and near-exact matches of individual instances. The term “exact matches” is self-explanatory; “near-exact matches” are completions by the LLM that, while not identical, show considerable overlap and maintain significant semantic and structural similarity to the reference instance. To generalize from individual instances to entire partitions, the human expert followed the following rule: a partition is flagged as contaminated if the instance-based evaluation identified at least one exact match or at least two near-exact matches.
@@
Evaluation Metrics: In our analysis, the computation of the BLEURT score varies based on the structure of the dataset/instance, as this metric hinges on the fluency and quality of the generated sequence. For single-instance datasets, where individual instances are randomly cut off mid-sentence and then completed by the LLM, we join the model-produced continuation to the severed reference instance and then calculate the BLEURT score. Conversely, for instances from paired-instance and multi-sentence single-instance datasets, the BLEURT score is computed solely for the newly produced sequence. We highlight that our BLEURT score computations use the most recent checkpoint provided, i.e., BLEURT-20 (Sellam, Das, and Parikh 2020; Pu et al. 2021; Devlin et al. 2019). On the other hand, regardless of the dataset/instance type, the ROUGE score calculation exclusively pertains to the portions of the text finished by the LLM. This is due to the score’s dependency on statistical attributes rather than semantic consistency.
Table 1: Overall accuracy at detecting contamination across 14 dataset partitions based on the human judgments. The two proposed algorithms are evaluated for GPT-4 (top half) and GPT-3.5 (bottom half).
Method Success Rate Accuracy
Base Model: GPT-4
Algorithm 1: BLEURT 11/14 78.57%
Algorithm 1: ROUGE-L 13/14 92.86%
Algorithm 2: GPT-4 ICL 14/14 100.00%
Base Model: GPT-3.5
Algorithm 1: BLEURT 9/14 64.29%
Algorithm 1: ROUGE-L 7/14 50.00%
Algorithm 2: GPT-4 ICL 13/14 92.86%
Table 2: An assessment of our proposed method in contrast to the comparative method, with the evaluations rooted in BLEURT and ROUGE-L scores, along with the decisions from GPT-4 and human judgments. The evaluations are performed on 10 instances randomly drawn from each split of a particular dataset, with GPT-4 (top half) and GPT-3.5 (bottom half) serving as the LLMs that are investigated. Asterisks indicate statistically significant differences between the guided and general instructions.
Metrics Instruct. Type IMDB AG News Yelp RTE WNLI SAMSum XSum
Base Model: GPT-4
BLEURT Score on Train General 0.43 0.63 0.43 0.54 0.47 0.58 0.54
@@
BLEURT Score on Test/Validation General 0.43 0.62 0.41 0.50 0.50 0.58 0.64
@@
ROUGE-L Score on Train General 0.14 0.17 0.15 0.41 0.26 0.13 0.18
@@
ROUGE-L Score on Test/Validation General 0.16 0.16 0.15 0.31 0.36 0.12 0.23
@@
GPT-4 Eval. on Train × ✓✓ × ×
GPT-4 Eval. on Test/Validation × ✓✓ × × ×
Did-ChatGPT-Cheat? on Train ? ? ? ? ? ? ?
Did-ChatGPT-Cheat? on Test/Validation ? ? ? ? ? ? ?
Human Eval. on Train × ✓✓ × ×
Human Eval. on Test/Validation × ✓✓ × × ×
Base Model: GPT-3.5
BLEURT Score on Train General 0.45 0.58 0.45 0.50 0.49 0.59 0.54
@@
BLEURT Score on Test/Validation General 0.45 0.60 0.42 0.47 0.47 0.58 0.62
@@
ROUGE-L Score on Train General 0.12 0.06 0.13 0.37 0.29 0.10 0.14
@@
ROUGE-L Score on Test/Validation General 0.13 0.10 0.11 0.23 0.32 0.13 0.18
@@
GPT-4 Eval. on Train × × × × × × ×
GPT-4 Eval. on Test/Validation × × × × × × ×
Did-ChatGPT-Cheat? on Train × × × × ? × ×
Did-ChatGPT-Cheat? on Test/Validation × × × × ? × ×
Human Eval. on Train × × × × × × ×
Human Eval. on Test/Validation × × × × × × ✓✓
Comparative Framework: The most relevant approach to detect data contamination in LLMs is Did-ChatGPT-Cheat? (Sainz et al. 2023), against which we contrast our proposed method. Unlike our method, which employs a binary scale to determine dataset contamination, the comparison approach includes a “suspicious” category. This designation is invoked when the LLM, upon being asked to generate the first instances of a specific partition in the dataset, outputs characteristic attributes such as data format, IDs, or other dataset-specific details instead of the actual initial instances. If the model, on the other hand, fails to produce these characteristics, it is deemed “not contaminated.”
5 Results and Discussion
Table 1 lists the overall accuracy of the proposed algorithms in 28 distinct settings: two language models (GPT-4 and GPT-3.5) $\times$ 14 dataset partitions coming from seven datasets. Table 2 provides a detailed breakdown of each algorithm per dataset partition and language model. Within this table, partition-level contamination is represented in the following ways: a single tick (✓) signifies the presence of at least one exact match, whereas a double tick (✓✓) indicates the detection of at least two near-exact matches. In contrast, the cross sign (×) denotes that neither of the aforementioned conditions were met. In the context of the Did-ChatGPT-Cheat? method, this cross sign indicates that the model’s output does not contain any specific information about the initial instances of the dataset partition upon the request to generate them. For the same method, the question mark (?) highlights partitions that are deemed suspicious. We draw the following observations from these experiments:
(1) Comparing the output of the guided instruction against that produced by the general instruction (Algorithm 1) proves to be an effective strategy for identifying data contamination in LLMs. The effectiveness of this strategy is highlighted by its ability to pinpoint either exact or near-exact matches in half of the evaluated settings (7 out of 14) when using GPT-4 as the underlying model. Moreover, in cases where the partition remains uncontaminated, the consistency between average BLEURT and ROUGE-L scores—drawn from both general and guided instructions—substantiates our assertion. This consistency suggests that establishing a baseline using random matches from general instruction is valid. Additionally, these outcomes further confirm that guided instruction effectively steers the LLM’s domain towards the respective dataset, achieving notably higher average BLEURT and ROUGE-L scores when the dataset partition exhibits contamination.
Further, we note that in the context of GPT-4, Algorithm 1 using the BLEURT score successfully labels data contamination in 11/14 settings, whereas ROUGE-L does so in 13/14 settings. However, when applied to GPT-3.5, Algorithm 1 struggles to differentiate between contaminated and clean partitions: BLEURT does so in 9/14 cases, while ROUGE-L manages in only 7/14 scenarios.
(2) Using Algorithm 2 and assessments with GPT-4 through ICL prompting, we find that this approach aligns closely with human evaluations. Specifically, in experiments run on GPT-4 and GPT-3.5, the match rates are 14/14 and 13/14, respectively. These accuracies are higher than any produced by Algorithm 1 and maintain consistency across the two language models.
(3) Upon assessing the results of our comparative approach, Did-ChatGPT-Cheat?, we discover that this method invariably labels partitions as “suspicious” for all scenarios involving GPT-4. This label provides no insight into the potential contamination of a split; hence, no division received the correct classification based on human evaluation (0/14). For GPT-3.5, based on the defined criteria for uncontaminated partitions, this method detects 11/13 settings correctly. This result supports our observation that detecting contamination at the instance level (before generalizing to the partition level) is a more robust strategy.
(4) Last but not least, the human evaluation reveals that the train and test/validation splits of both the AG News and WNLI datasets were included in GPT-4’s training data. However, for IMDB and RTE, only the training portions were incorporated, while for XSum, only the test split was leaked. For GPT-3.5, the only data exposure was the test partition of the XSum dataset. This finding confirms that, despite their creators’ efforts, today’s LLMs have ingested NLP datasets. We hope that this observation will inform the design of better scientific experiments in the NLP space in the future.

### v2

@@ 3.2 Detecting Partition-level Contamination
To generalize from instance-level contamination to partition-level discrete decisions (i.e., the partition is/is not contaminated), we take advantage of two observations:
@@
Idea 1: A dataset is likely to be contaminated if the average overlap score with the reference instances (as measured by ROUGE-L and BLEURT) observed with completions from the guided instruction is significantly larger than the one measured with the completions from the general instruction. The motivation behind this idea is that since the only difference between the two instructions is that the guided instruction contains the dataset and partition name as guidance, the improvement can only be explained by contamination.
Idea 2: A dataset is likely to be contaminated if GPT-4 using few-shot ICL prompting detects at least one exact match or at least two near-exact matches. The intuition behind this idea is that even a small contaminated part of the sample of instances is likely indicative of a larger dataset partition leak. While the presence of an exact match among replicas generated by LLM is a clear sign of contamination, the approach to handling exact or near-exact matches—and deciding the number of such matches that indicates broader contamination—can be tailored depending on specific research objectives. In this paper, we intuitively establish the above-mentioned criterion to extrapolate from the instance-level to the partition-level contamination. An empirical validation of our approach is also provided in Section 3.3.
We propose two algorithms, each implementing one of these ideas respectively.
@@
Algorithm 1: A dataset partition is labeled as contaminated if the average overlap score (as provided by BLEURT and ROUGE-L) between the reference instances and generated texts with the guided instruction on a sample of ten instances is statistically significantly better than those produced by general instruction under a non-parametric bootstrap resampling test. [^2] Details of our bootstrap resampling method can be found in Appendix D.
@@
The advantage of this algorithm is that it is non-parametric, i.e., we do not need to decide on an arbitrary threshold on the ROUGE-L or BLEURT scores to indicate contamination. However, its drawback is that even a significant increase in overlap may still come from generated instances that a human would not consider an exact or near-exact match. Algorithm 2 addresses this limitation.
@@
Algorithm 2: A dataset partition is labeled as contaminated if GPT-4 with few-shot ICL prompting flags at least one generated completion as an exact match or a minimum of two completions as near-exact matches within a sample of ten instances. All completions in this setting are generated solely by guided instruction.
We evaluate both these algorithms in Section 5.
3.3 Instance Replication: A Valid Approach to Detect Data Contamination
To validate our choice for the hyperparameters used in Algorithm 2, i.e., the number of exact/near-exact matches to declare contamination, we performed a controlled study in which an LLM is contaminated on purpose with several datasets. To this end, we used the GPT-3.5 base model and a subset of the train partition of the following datasets (one dataset from each task in question): AG News, RTE, and XSum. Note that all these partitions were marked as uncontaminated for GPT-3.5 by the human evaluators (see Table 4 and Section 4: Human Evaluation). To mimic the LLM’s pre-training on web data, we retained only minimal metadata about the datasets as they appear on the web when scraped. In particular, we used: the dataset title, the partition name, and the entire instance. [^3] All data formats used for the contamination of GPT-3.5 are detailed in Table 9 in Appendix E. Following training, we evaluate the generated completions by our best-performing technique (Algorithm 2: GPT-4 ICL) (see Table 3). Figure 2 visualizes the generated replicas before and after contamination in one of our experiments when guided instruction is utilized. [^4] Further examples are provided in Table 10 in Appendix F. In addition, Table 2 summarizes our findings from this study. The key conclusion of this experiment is that the contaminated LLM generated at least one exact match in each setting. This underscores that the replication of even one exact match stands as a robust and undeniable indicator of contamination. [^5] Details on the continued training of the GPT-3.5 base model is presented in Appendix E.
As a second experiment, we employed GPT-4 and the GSM8k dataset (Cobbe et al. 2021). This choice was motivated by OpenAI’s technical report on GPT-4, which indicates contamination from its train split (OpenAI 2023). Given that this dataset comprises mathematical problems, our objective is to replicate the questions in the dataset while withholding their corresponding answers. [^6] An example of this replication process is provided in Table 10 in Appendix F. Table 2 reports our results from this experiment. Our results highlight that contamination is not solely identified through exact matches; near-exact matches are also indicative. To account for the probabilistic nature of LLMs, we set a threshold of two for the minimum number of near-exact matches to indicate contamination. As shown, this is supported by the data.
Instruction: You are provided with Sentence 1 from the train split of the RTE dataset. Finish Sentence 2 as appeared in the dataset. Sentence 2 must exactly match the instance in the dataset. Sentence 1: Twelve of Jupiter’s moons are relatively small and seem to have been more likely captured than to have been formed in orbit around Jupiter. Label: 0 (not entailment) Sentence 2: ——————————————————————————————————————————————— GPT-3.5: The formation of Jupiter’s twelve relatively small moons is more likely due to their capture rather than being formed in orbit around Jupiter. ——————————————————————————————————————————————— Contaminated GPT-3.5: Jupiter has twelve moons.
Figure 2: An example of an exact match generated by the GPT-3.5 contaminated with the train split of the RTE dataset versus an inexact match generated by the GPT-3.5 base model, both under the same guided instruction. This example is one of the training instances used during contamination.
Table 1: Results after introducing intentional contamination to the GPT-3.5 base model using guided instruction. A tick (✓) indicates the identification of at least one exact replica from the training instances used for contamination by our top-performing method (Alg. 2: GPT-4 ICL) and human evaluation.
Method AG News RTE XSum
Alg. 2: GPT-4 ICL
Human Evaluation
Table 2: Results of identifying contamination of GSM8k dataset within GPT-4 when guided instruction is used. A double tick (✓✓) signals the identification of two or more near-exact replicas from the train split of this dataset by our top-performing method (Alg. 2: GPT-4 ICL) and human evaluation.
Method GSM8k
Alg. 2: GPT-4 ICL ✓✓
Human Evaluation ✓✓
@@ 4 Experimental Setup
Data: Our evaluation employs seven datasets derived from various tasks, namely classification, summarization, and NLI. The datasets in question involve IMDB (Maas et al. 2011), AG News (Zhang et al. 2015), Yelp Full Reviews (Zhang et al. 2015), SAMSum (Gliwa et al. 2019), XSum (Narayan et al. 2018), WNLI (Wang et al. 2018), and RTE (Wang et al. 2019). In order to ensure a comprehensive experimental setup, all our experiments are carried out on both the training and test/validation splits of the aforesaid datasets. We make use of the publicly available divisions, working with the training and test splits for each. However, for the last two datasets, only the validation splits were publicly accessible with their labels. Considering our research’s emphasis on pinpointing data contamination with minimal dataset instances, the resource constraints, and our intention to facilitate the replication of this approach by other researchers, we randomly chose 10 instances from each split for our experiments.
@@
Setting: We use snapshots of GPT-3.5 and GPT-4 from June 13, 2023—specifically gpt-3.5-turbo-0613 and gpt-4-0613—both accessed via the OpenAI API, as our foundation LLMs. To obtain deterministic results, we set the temperature to zero and capped the maximum completion length at 500 tokens. Contrarily, our comparative method (ChatGPT-Cheat?) uses the chat user interface (UI), which we also leveraged for conducting the experiment under this method. Specifically, we used the UI versions of GPT-4 and GPT-3.5 that were released on July 20, 2023.
@@
Human Evaluation: We undertake a human evaluation, led by two domain experts, [^7] The two annotators had almost perfect inter-rater agreement across all settings. This is due to the fact that a small subset of instances was used for contamination detection, and contamination is evident when it occurs. to characterize contamination by identifying both exact matches and near-exact matches of individual instances. The term “exact matches” is self-explanatory; “near-exact matches” are completions by the LLM that, while not identical, show considerable overlap and maintain significant semantic and structural similarity to the reference instance. To generalize from individual instances to entire partitions, the human annotators followed the rule described in Algorithm 2 that was validated empirically in Section 3.3: a partition is flagged as contaminated if the instance-based evaluation identifies at least one exact match or at least two near-exact matches.
@@
Evaluation Metrics: In our analysis, the computation of the BLEURT score varies based on the structure of the dataset/instance, as this metric hinges on the fluency and quality of the generated sequence. For single-instance datasets, where individual instances are randomly cut off mid-sentence and then completed by the LLM, we join the model-produced continuation to the severed reference instance and then calculate the BLEURT score. Conversely, for instances from paired-instance and multi-sentence single-instance datasets, the BLEURT score is computed solely for the newly produced sequence. We highlight that our BLEURT score computations use the most recent checkpoint provided, i.e., BLEURT-20 (Pu et al. 2021). On the other hand, regardless of the dataset/instance type, the ROUGE-L score calculation exclusively pertains to the portions of the text finished by the LLM. This is due to the score’s dependency on statistical attributes rather than semantic consistency.
Comparative Framework: We compare our proposed methods against the ChatGPT-Cheat? method (Sainz et al. 2023). Unlike our method, which uses a binary scale to determine contamination, the comparison approach includes a “suspicious” category. This designation is invoked when the LLM, upon being asked to generate the first instances of a dataset split, outputs characteristic attributes such as data format, IDs, or other dataset-specific details instead of the actual instances. If the model, on the other hand, fails to produce these characteristics, it is deemed uncontaminated.
5 Results and Discussion
Table 3 lists the overall accuracy of our proposed methods in 28 distinct settings: two LLMs (GPT-4 and GPT-3.5) $\times$ 14 dataset partitions coming from seven datasets. Table 4 provides a detailed breakdown of each method per dataset partition and the respective LLM. We draw the following observations from our experiments:
Table 3: Overall accuracy at detecting contamination across 14 partitions for GPT-4 and GPT-3.5. The two LLMs are evaluated against human annotators. The “Success Rate” shows how often each method matches human judgment, while the “Accuracy” gives the corresponding percentages.
GPT-4 GPT-3.5
Method Success Rate Accuracy Success Rate Accuracy
Strict Eval.: ChatGPT-Cheat? 0/14 0.00% 11/14 78.57%
Lenient Eval.: ChatGPT-Cheat? 9/14 64.29% 13/14 92.86%
Algorithm 1: BLEURT 11/14 78.57% 9/14 64.29%
Algorithm 1: ROUGE-L 13/14 92.86% 7/14 50.00%
Algorithm 2: GPT-4 ICL 14/14 100.00% 13/14 92.86%
(1) Algorithm 1, which hinges on the difference in average overlap scores between outputs from guided instruction and those from general instruction, performs well in the majority of settings. Its best performance is a success rate of 13/14 when using GPT-4 as the underlying model and 9/14 when using GPT-3.5. We consider these results exciting given the algorithm’s simplicity. However, Table 3 shows that: (a) its performance is not universally good—it performs at chance level when using ROUGE-L on GPT-3.5 outputs (7/14), and (b) its success rate varies depending on the metric in use (i.e., BLEURT or ROUGE-L).
(2) In contrast, Algorithm 2, which relies on GPT-4 evaluation using the few-shot ICL prompt, aligns closely with human evaluations. Specifically, in experiments run on GPT-4 and GPT-3.5, its success rates are 14/14 and 13/14, respectively. These accuracies are higher than any produced by Algorithm 1 and maintain consistency across all the settings with the two LLMs.
(3) Upon assessing the results of ChatGPT-Cheat? method, we discover that this method invariably labels partitions as suspicious—likely due to the precaution against generating copyrighted content which is activated by safety filters—for all scenarios involving GPT-4. Given this, we interpret the outcomes of this method through two lenses: strict and lenient evaluation. In the strict evaluation, we do not interpret the suspicious label as contaminated or uncontaminated. Under this assessment, no partition is correctly classified according to human evaluation (0/14) in settings with GPT-4, and 11/14 in settings with GPT-3.5. In the lenient evaluation, we convert the suspicious label to either contaminated or uncontaminated in a way that maximizes the performance of this method. In this setting, the ChatGPT-Cheat? method correctly identifies 9/14 and 13/14 in settings with GPT-4 and GPT-3.5, respectively. However, this lenient evaluation is unrealistic due to the overfitting in interpreting the suspicious label. These findings support our observation that identifying contamination at the instance level, before extrapolating to the partition level, is a more resilient strategy.
(4) Last but not least, the human evaluation reveals that the train and test/validation splits of both the AG News and WNLI datasets were included in GPT-4’s pre-training data. However, for IMDB and RTE, only the training partitions were incorporated, while for XSum, only the test split was leaked. For GPT-3.5, the only data exposure was the test partition of the XSum dataset. These findings confirm that, despite their creators’ efforts, today’s LLMs have ingested NLP datasets. We hope that this observation informs the design of better scientific experiments with LLMs in the NLP space in the future.
Table 4: An assessment of our proposed methods in contrast to ChatGPT-Cheat? method. We evaluate Algorithm 1 using BLEURT and ROUGE-L, as well as Algorithm 2 which relies on GPT-4 decisions via few-shot ICL prompting. The evaluations are performed on 10 instances randomly drawn from each split of a particular dataset, with GPT-4 and GPT-3.5 serving as the LLMs that are investigated. Partition-level contamination is represented in the following ways: (1) While asterisks (*) indicate statistically significant differences between the completions produced by guided and general instructions (as measured by BLEURT and ROUGE-L), underlined numbers indicate settings that align with human evaluations (Algorithm 1). (2) A single tick (✓) points to the presence of at least one exact match, while a double tick (✓✓) signals the identification of two or more near-exact matches (Algorithm 2). A cross sign (×) denotes that neither of the aforementioned conditions were met. For the ChatGPT-Cheat? method, this cross sign indicates that the model’s output does not contain any specific information about the first instances of the dataset partition upon the request to generate them. For the same method, the question mark (?) highlights partitions that are deemed suspicious.
Model Method Split Instruct. Datasets
IMDB AG News Yelp RTE WNLI SAMSum XSum
GPT-4 Alg. 1: BLEURT Train General 0.43 0.63 0.43 0.54 0.47 0.58 0.54
@@ 5 Results and Discussion
Test/Valid General 0.43 0.62 0.41 0.50 0.50 0.58 0.64
@@
Alg. 1: ROUGE-L Train General 0.14 0.17 0.15 0.41 0.26 0.13 0.18
@@
Test/Valid General 0.16 0.16 0.15 0.31 0.36 0.12 0.23
@@
Alg. 2: GPT-4 ICL Train Guided × ✓✓ × ×
Test/Valid Guided × ✓✓ × × ×
ChatGPT-Cheat? Train Guided ? ? ? ? ? ? ?
Test/Valid Guided ? ? ? ? ? ? ?
Human Evaluation Train Guided × ✓✓ × ×
Test/Valid Guided × ✓✓ × × ×
GPT-3.5 Alg. 1: BLEURT Train General 0.45 0.58 0.45 0.50 0.49 0.59 0.54
@@
Test/Valid General 0.45 0.60 0.42 0.47 0.47 0.58 0.62
@@
Alg. 1: ROUGE-L Train General 0.12 0.06 0.13 0.37 0.29 0.10 0.14
@@
Test/Valid General 0.13 0.10 0.11 0.23 0.32 0.13 0.18
@@
Alg. 2: GPT-4 ICL Train Guided × × × × × × ×
Test/Valid Guided × × × × × × ×
ChatGPT-Cheat? Train Guided × × × × ? × ×
Test/Valid Guided × × × × ? × ×
Human Evaluation Train Guided × × × × × × ×
Test/Valid Guided × × × × × × ✓✓